# Deep learning-derived arterial input function

### v1

@@ 3.2 Basis functions
Given that AIFs generally exhibit certain smoothness characteristics, such as typically smoother tails, employing basis functions for their parameterization is beneficial for maintaining this smoothness in the estimations. An additional benefit of employing basis functions, as opposed to direct AIF estimation, is their ability to provide continuous functions. Direct estimation, in contrast, can only yield values at discrete time points and may require potentially error-prone interpolation to fill in values between these points. In our study, we explored the use of Gaussian and truncated exponential functions as basis functions. These have historically been used for modeling AIFs due to their inherent properties that align well with the characteristics of AIFs. Gaussian functions, for instance, have been employed in [37, 45], while exponential functions have seen use in [17, 46]. However, unlike these traditional methods where the number, scale, and sign of the basis functions are predetermined, our method capitalizes on the complex modeling capabilities of DNNs. This allows the DNN to freely learn and adjust all parameters of the basis functions to achieve the most accurate fit.
@@ 3.2.2 Exponential-sigmoid Basis Function
When using exponential functions for modeling AIF, it is important to consider that these functions do not inherently reduce to zero. Yet, the AIF value prior to the injection of the radioactive tracer is in fact zero. To accommodate this, previous methods have employed a truncated exponential model that introduces a discontinuity for the AIF [17, 46]. In such models, the estimated IDIF is set to zero before a specified time point. After this time point, the IDIF is then modeled to follow a form of exponential functions, aligning with the behavior observed pre- and post-tracer injection. However, the discontinuity inherent in previous AIF models using exponential functions poses a challenge, as it is not differentiable and thus prevents the backpropagation of gradients needed to update DNN parameters during training. To address this, we have devised a workaround by relaxing the discontinuity with a sigmoid function, as similarly done in by [45]. In our model, the superposition of $K$ exponential-sigmoid functions is formulated as:
@@ 3.3 Scaling factor estimation head
This echoes the approach commonly used in traditional decomposition-based IDIF estimation techniques [40, 41, 3], where blood sampling at the peak of AIF is used to scale the IDIF. However, in our model, the scaling factor is determined by the DNN. Through empirical evaluation, we have found that this approach of using a scaling factor yields better results than having the DNN directly output the IDIF without any scaling. The evidence supporting this finding is detailed in a later section of the paper.
@@ 4.1 Dataset and preprocessing
The dataset used in this work were de-identified dynamic 11C-DPA-713 (DPA) brain PET images. DPA is a second generation PET tracer targeting the translocator protein 18kDa (TSPO) for detecting microglial response or proliferation in vivo  [5, 59, 57, 39, 9, 62]. It has shown superior binding affinity and signal-to-background ratio [59, 57, 39, 9, 62]. The DPA affiliation is affected by single nucleotide polymorphism (rs6971) TSPO genotyping, with C/C, C/T and T/T coresponding to high-affinity binders (HAB), mixed-affinity binders (MAB), and low-affinity binders (LAB), respectively.[43, 36] The data used in this studies were collected through several clinical research studies at Johns Hopkins University that used DPA PET in health control individuals [16], and several conditions that may cause neuroinflammation, such as HIV, Lyme disease, and repeated traumatic brain injury [10, 11, 52, 50, 51].
From these studies, de-identified data from HAB and MAB participants were collected, including 37 controls and 86 patients. Among those 67 were HAB and 56 were MAB. The data includes 90 minutes dynamic PET images and metabolite corrected AIF measured from blood sampling. The PET data were acquired on a brain-dedicated High Resolution Research Tomograph (HRRT, Siemens Healthcare, Knoxville, TN), with a fitted thermoplastic facemask for head fixation to reduce motion. The 90 min data were binned into 30 frames and reconstructed using the iterative ordered subsets expectation maximization algorithm [11]. The data were acquired from each participant through studies approved by the Johns Hopkins Institutional Review Board. Each participant provided written informed consent, which included the use of their data in secondary analyses.
All dynamic PET images of the brain were first aligned with a template through affine registration. Subsequently, skull stripping was applied to exclude non-brain regions. This procedure was carried out using SynthStrip, a publicly available learning-based method [24]. In particular, SynthStrip was applied to an image averaged over all time frames, and the generated mask was then used to consistently strip the brain regions across the time frames. Subsequently, the images were uniformly cropped to a size of $160\times 192\times 160$, maintaining an isotropic resolution of 1 $mm$. Additionally, the uptakes in the images were normalized to the Standardized Uptake Value (SUV), and the input function was correspondingly normalized to ensure consistency in the data. For training and evaluation, we implemented five-fold cross-validation, with each fold comprising an 8:2 split for training and testing.
@@ 4.2 Implementation details
The models were implemented using the PyTorch framework [47] on a PC equipped with two NVIDIA Quadro P6000 GPUs and an NVIDIA RTX A4000 GPU. The training was conducted over 500 epochs, using the Adam optimizer [27] with a learning rate of 1e-4 and a batch size of 1. To enhance the robustness of the models, the dataset underwent augmentation through random spatial direction flipping during the training phase.
@@ 4.3 Baseline methods
To evaluate the effectiveness of our proposed method, we conducted comparisons with two popular convolutional neural networks, ResNet50 [23] and ConvNeXt [34], as well as a traditional IDIF estimation method, EPICA [40, 41], that predicts AIF directly from dynamic PET image sequences. The specifics of these baseline methods are outlined below.
  • ResNet50.—We first compare the proposed network architecture with a widely used ConvNet architecture, ResNet50 [23], which has found extensive applications across a diverse range of image classification and regression tasks. To ensure a fair comparison, the input configuration for ResNet50 mirrors that of our proposed network. This involves reducing the original spatial resolution of the dynamic PET sequence by a factor of 4 and concatenating it along the temporal axis, resulting in the formation of $T$ channels.
  • ConvNeXt.— We then evaluate the proposed DLIF against ConvNeXt [34], a recent ConvNet architecture that has been recognized as a strong counterpart to Transformer models. We employ the default configuration of ConvNeXt but convert its 2D modules into 3D modules with the input remaining consistent with the aforementioned specifications.
  • EPICA.—The final baseline method compared in this study is EPICA [40, 41], a statistical decomposition-based method discussed in the Introduction section. EPICA requires a scaling factor, typically derived from a single blood sample, to calibrate the estimated IDIF to align with the amplitude of the actual AIF. In our evaluation, we used the peak value of the ground-truth AIF as the scaling factor for the estimated IDIF.
@@ 4.4 Evaluation metrics
The DLIFs generated by the proposed framework have been applied in Logan graphical analysis. However, it is important to note that the potential applications of the proposed DLIF are not confined to this specific analysis technique. Logan graphical analysis employs linear regression to analyze the data after a specified time, estimating the slope of the resulting line to determine the total volume of distribution, as discussed in [4]. In this context, an exact match between the DLIF and the actual AIF before this chosen time is not critical. The crucial aspect is that the integral of the DLIF (i.e., the area under its curve) matches that of the AIF prior to this time. After this time point, the curves should align as closely as possible. This precise alignment is vital for accurately estimating the slope, which is crucial for the correct calculation of the total volume of distribution. Given these considerations, using metrics such as the mean squared error or the mean absolute error to assess the accuracy of the DLIFs may not provide the most meaningful insights. This is because such metrics could be disproportionately influenced by the errors in matching the peak values, which are significantly higher in amplitude compared to the input function values at later time points.
4.4.1 Pearson’s correlation coefficient ($\boldsymbol{r}$
To evaluate the accuracy of AIFs estimated by various methods, we adopt Pearson’s correlation coefficient or $\boldsymbol{r}$, a metric also employed in related research on AIF estimation in perfusion CT [49]. This metric quantifies the correlation between the estimated AIF and the ground truth, and is mathematically defined as:
@@ 4.4.4 Logan graphical analysis
The Logan graphical analysis is widely used for quantifying dynamic PET images of a reversible tracer, where the tissue time-activity curve (TAC) is mathematically transformed and plotted against "normalized time" to estimate the total distribution volume ($V_{T}$). It is mathematically defined as:
[…]
, where $C_{T}(t)$ is the measured tissue activity and $C_{p}(t)$ is the input curve. For the reversible compartment, this expression results in a straight line with a slope of $V_{T}$ after an equilibration time (t*). In previous studies, we have demonstrated that Logan graphical analysis with a t* = 30 minutes could provide robust $V_{T}$ values that agree with those produced from compartmental models [16].
@@ 4.5 Ablation study and component analysis
  • "Direct Est."—This model directly predicts AIF values at specific discrete intervals, in accordance with the AIF values acquired using arterial sampling.
  • "Exp."—This model generates parameters for the exponential-sigmoid basis functions, which are then superimposed to create a continuous estimate of the AIF.
  • "Exp. + Peak"—In addition to generating parameters for the exponential-sigmoid basis functions, this model adds a scaling factor to adjust the amplitude of the estimated AIF.
  • "Exp. + Sparse"—In addition to generating parameters for the exponential-sigmoid basis functions, this model applies a sparsity constraint to encourage employing a smaller set of basis functions.
  • "Exp. + Peak + Sparse"—This model not only generates parameters for the exponential-sigmoid basis functions and a scaling factor but also includes a sparsity constraint.
  • "Gaussian"—This model outputs parameters for Gaussian basis functions, which are used to model the AIF as a sum of Gaussian curves.
  • "Gaussian + Peak"—In addition to the Gaussian basis function parameters, this model includes a scaling factor to scale the amplitude of the estimated AIF.
  • "Gaussian + Sparse"—In addition to the Gaussian basis function parameters, this model adds a sparsity constraint on the Gaussian basis functions, aiming to reduce the number of basis functions required for representing the AIF.
  • "Gaussian + Peak + Sparse"—This model generates parameters for the Gaussian basis functions, includes a peak scaling factor and applies a sparsity constraint during training.
Fig. 6: Qualitative comparison of baseline versus proposed techniques using five-fold cross-validation on both genotypes. The first two columns present results from patients with HAB, contrasted with the last two columns showcasing results from MAB patients. The first row depicts overall estimated input functions, and the second and the third rows offer zoom-in views within the 0 to 7 minute and 22 to 90 minute timeframes, respectively.
@@ 5.1 Composition of DLIFs using basis functions
First, configurations employing Gaussian basis functions typically activate a larger number of functions compared to those using exponential-sigmoid basis functions. Even when applying a sparsity constraint, multiple Gaussian functions remain active (e.g., "Gaussian + Peak + Sparse"), enabling DLIFs to accurately capture subtle details of the ground-truth AIF. Conversely, exponential-sigmoid configurations often utilize fewer basis functions. Notably, under sparsity constraints, the DNN can converge to solutions involving a minimal number of exponential-sigmoid functions—such as the "Exp. + Peak + Sparse" configuration—where only one function remains active, effectively suppressing the others. This outcome highlights the intrinsic compatibility between the exponential-sigmoid shape and true AIF curves, facilitating concise yet accurate representations.
@@ 5.2 Ablation and component analysis
Next, we analyze the key elements that contribute to the improved quantitative performance of the proposed DLIF framework. To evaluate this, we used Pearson’s correlation coefficient ($\boldsymbol{r}$) to quantify the alignment of the DLIF shapes with the ground-truth AIFs. Additionally, we introduce a novel metric, Intersection-over-Union (IoU), which measures the overlap between the areas under the estimated and true input function curves. Unlike metrics focused on matching peak values, IoU emphasizes the integral areas, which are critical for downstream analyses such as Logan graphical analysis [35]. Details of these metrics are provided in the Methods section.
Table 1: Quantitative analysis of different methods on test sets not exposed during training. The results are an average from five-fold cross-validation. Note that the time intervals "30-" and "30+" refer to the periods of 0-30 minutes and 60-90 minutes, respectively. The top-performing scores are highlighted in bold. The symbol "*" denotes a statistically significant improvement, with a $p$-value < 0.05, as determined by a Wilcoxon signed-rank test with Bonferroni adjustment for multiple comparisons.
@@
  EPICA 0.50$\pm$0.29 0.39$\pm$0.16 0.32$\pm$0.26 2.34$\pm$1.22 0.23$\pm$0.23 - 0.52$\pm$0.27 0.36$\pm$0.13 0.34$\pm$0.28 3.48$\pm$1.14 0.25$\pm$0.12 -
ResNet50 0.90$\pm$0.11 0.62$\pm$0.15 0.55$\pm$0.14 1.72$\pm$1.10 0.14$\pm$0.10 0.52$\pm$0.52 0.90$\pm$0.09 0.57$\pm$0.17 0.56$\pm$0.15 2.36$\pm$1.15 0.14$\pm$0.07 0.45$\pm$0.20
ConvNeXt 0.93$\pm$0.09 0.70$\pm$0.13 0.74$\pm$0.12 1.25$\pm$0.66 0.07$\pm$0.03 0.37$\pm$0.31 0.94$\pm$0.05 0.73$\pm$0.10 0.73$\pm$0.11 1.66$\pm$0.85 0.09$\pm$0.05 0.28$\pm$0.16
DLIF-Exp. 0.94$\pm$0.08 0.72$\pm$0.13 0.73$\pm$0.14 1.13$\pm$0.66* 0.07$\pm$0.04 0.33$\pm$0.26 0.94$\pm$0.07 0.73$\pm$0.13 0.74$\pm$0.13 1.65$\pm$1.07 0.08$\pm$0.05 0.27$\pm$0.17
DLIF-Gaussian+Peak 0.94$\pm$0.08 0.71$\pm$0.13 0.76$\pm$0.14* 1.19$\pm$0.75 0.06$\pm$0.05 0.35$\pm$0.34 0.94$\pm$0.07 0.72$\pm$0.11 0.77$\pm$0.13* 1.70$\pm$0.91 0.07$\pm$0.05* 0.27$\pm$0.17
DLIF-Gaussian+Peak+Sparse 0.93$\pm$0.09 0.72$\pm$0.13 0.75$\pm$0.13 1.17$\pm$0.74 0.06$\pm$0.04* 0.31$\pm$0.31* 0.94$\pm$0.07 0.72$\pm$0.13 0.77$\pm$0.15 1.67$\pm$1.01 0.07$\pm$0.05* 0.26$\pm$0.17
To refine the evaluation, we partition the input functions into two temporal segments: before and after 30 minutes, corresponding to the peak and stable phases of the input functions. Figure 5 presents the results of the component analysis using five-fold cross-validation on the validation datasets. The first plot in Fig. 5 depicts $\boldsymbol{r}$ values, representing the shape correspondence between the DLIFs and the actual AIFs. Although direct estimation of input functions (labeled "Direct Est.") achieves an average $\boldsymbol{r}$ of 0.958, DLIF configurations using basis functions, either exponential or Gaussian, consistently outperform this approach. Specifically, the "Exp.", "Gaussian", "Gaussian+Peak", and "Gaussian+Peak+Sparse" configurations attain average $\boldsymbol{r}$ values of 0.965, 0.964, 0.961, and 0.964, respectively. The second plot evaluates IoU, capturing the agreement in amplitudes and the total areas under the curves. Here, the "Exp.", "Gaussian", "Gaussian+Peak", and "Gaussian+Peak+Sparse" configurations achieve mean IoU scores of 0.787, 0.777, 0.781, and 0.780, respectively, slightly outperforming "Direct Est." (0.776).
Performance in the different temporal phases reveals further insight. During the peak phase (<30 minutes), the top-performing configurations, "Exp.", "Gaussian+Peak", and "Gaussian+Peak+Sparse", record average IoU scores of 0.786, 0.781, and 0.779, respectively, compared to 0.776 for "Direct Est.". During the tail phase (>30 minutes), these configurations achieve average IoU scores of 0.808, 0.803, and 0.799, respectively, compared to 0.780 for "Direct Est.". Notably, the "Gaussian" configuration slightly underperforms in the peak phase, with an average IoU of 0.776.
Furthermore, the scatter plots demonstrate tighter distributions with shorter tails for "Exp.", "Gaussian+Peak", and "Gaussian+Peak+Sparse" compared to "Direct Est.", indicating more consistent performance. Based on these findings, we selected the top three configurations—"Exp.", "Gaussian+Peak", and "Gaussian+Peak+Sparse"—–for further quantitative and qualitative comparisons against baseline methods, as detailed in the subsequent section.

### v2

@@ 3.2 Basis functions
Given that AIFs generally exhibit certain smoothness characteristics, such as typically smoother tails, employing basis functions for their parameterization is beneficial for maintaining this smoothness in the estimations. An additional benefit of employing basis functions, as opposed to direct AIF estimation, is their ability to provide continuous functions. Direct estimation, in contrast, can only yield values at discrete time points and may require potentially error-prone interpolation to fill in values between these points. In our study, we explored the use of Gaussian and truncated exponential functions as basis functions. These have historically been used for modeling AIFs due to their inherent properties that align well with the characteristics of AIFs. Gaussian functions, for instance, have been employed in [mlynash2005automated, parker2006experimentally], while exponential functions have seen use in [feng1993models, parsey2000validation]. However, unlike these traditional methods where the number, scale, and sign of the basis functions are predetermined, our method capitalizes on the complex modeling capabilities of DNNs. This allows the DNN to freely learn and adjust all parameters of the basis functions to achieve the most accurate fit.
@@ 3.2.2 Exponential-sigmoid Basis Function
When using exponential functions for modeling AIF, it is important to consider that these functions do not inherently reduce to zero. Yet, the AIF value prior to the injection of the radioactive tracer is in fact zero. To accommodate this, previous methods have employed a truncated exponential model that introduces a discontinuity for the AIF [feng1993models, parsey2000validation]. In such models, the estimated IDIF is set to zero before a specified time point. After this time point, the IDIF is then modeled to follow a form of exponential functions, aligning with the behavior observed pre- and post-tracer injection. However, the discontinuity inherent in previous AIF models using exponential functions poses a challenge, as it is not differentiable and thus prevents the backpropagation of gradients needed to update DNN parameters during training. To address this, we have devised a workaround by relaxing the discontinuity with a sigmoid function, as similarly done in by [parker2006experimentally]. In our model, the superposition of $K$ exponential-sigmoid functions is formulated as:
@@ 3.3 Scaling factor estimation head
This echoes the approach commonly used in traditional decomposition-based IDIF estimation techniques [naganawa2005extraction, naganawa2005omission, b2006extraction], where blood sampling at the peak of AIF is used to scale the IDIF. However, in our model, the scaling factor is determined by the DNN. Through empirical evaluation, we have found that this approach of using a scaling factor yields better results than having the DNN directly output the IDIF without any scaling. The evidence supporting this finding is detailed in a later section of the paper.
@@ 4.1 Dataset and preprocessing
The dataset used in this work were de-identified dynamic [11C]DPA-713 brain PET images. [11C]DPA-713 is a second generation PET tracer targeting the translocator protein 18kDa (TSPO) for detecting microglial response or proliferation in vivo  [Chauveau_2008, Venneti_2006, Tichauer_2015, Muzi_2012, coughlin2014regional, Wang_2016]. It has shown superior binding affinity and signal-to-background ratio [Venneti_2006, Tichauer_2015, Muzi_2012, coughlin2014regional, Wang_2016]. The [11C]DPA-713 affiliation is affected by single nucleotide polymorphism (rs6971) TSPO genotyping, with C/C, C/T and T/T corresponding to high-affinity binders (HAB), mixed-affinity binders (MAB), and low-affinity binders (LAB), respectively.[Owen201118kDa, Milenkovic2018Effects] The data used in this study were collected through several clinical research studies at Johns Hopkins University that used [11C]DPA-713-TSPO-PET in health control individuals [endres2009initial], and several conditions that may cause neuroinflammation, such as HIV, Lyme disease, and repeated traumatic brain injury [Coughlin_2015, Coughlin_2018, rubin2018microglial, rubin2023imaging, rubin2022imaging].
From these studies, de-identified data from HAB and MAB participants were collected, including 37 controls and 86 patients. Among those 67 were HAB and 56 were MAB. The data includes 90 minutes dynamic PET images and metabolite corrected AIF measured from blood sampling. The PET data were acquired on a brain-dedicated High Resolution Research Tomograph (HRRT, Siemens Healthcare, Knoxville, TN), with a fitted thermoplastic facemask for head fixation to reduce motion. The 90 min data were binned into 30 frames and reconstructed using the iterative ordered subsets expectation maximization algorithm [Coughlin_2018]. The data were acquired from each participant through studies approved by the Johns Hopkins Institutional Review Board. Each participant provided written informed consent, which included the use of their data in secondary analyses.
All dynamic PET images of the brain were first aligned with a template through affine registration. Subsequently, skull stripping was applied to exclude non-brain regions. This procedure was carried out using SynthStrip, a publicly available learning-based method [hoopes2022synthstrip]. In particular, SynthStrip was applied to an image averaged over all time frames, and the generated mask was then used to consistently strip the brain regions across the time frames. Subsequently, the images were uniformly cropped to a size of $160\times 192\times 160$, maintaining an isotropic resolution of 1 $mm$. Additionally, the uptakes in the images were normalized to the Standardized Uptake Value (SUV), and the input function was correspondingly normalized to ensure consistency in the data. For training and evaluation, we implemented five-fold cross-validation, with each fold comprising an 8:2 split for training and testing.
@@ 4.2 Implementation details
The models were implemented using the PyTorch framework [paszke2019pytorch] on a PC equipped with two NVIDIA Quadro P6000 GPUs and an NVIDIA RTX A4000 GPU. The training was conducted over 500 epochs, using the Adam optimizer [kingma2014adam] with a learning rate of 1e-4 and a batch size of 1. To enhance the robustness of the models, the dataset underwent augmentation through random spatial direction flipping during the training phase.
@@ 4.3 Baseline methods
To evaluate the effectiveness of our proposed method, we conducted comparisons with two popular convolutional neural networks, ResNet50 [he2016deep] and ConvNeXt [liu2022convnet], as well as a traditional IDIF estimation method, EPICA [naganawa2005extraction, naganawa2005omission], that predicts AIF directly from dynamic PET image sequences. The specifics of these baseline methods are outlined below.
  • ResNet50—We first compare the proposed network architecture with a widely used ConvNet architecture, ResNet50 [he2016deep], which has found extensive applications across a diverse range of image classification and regression tasks. To ensure a fair comparison, the input configuration for ResNet50 mirrors that of our proposed network. This involves reducing the original spatial resolution of the dynamic PET sequence by a factor of 4 and concatenating it along the temporal axis, resulting in the formation of $T$ channels.
  • ConvNeXt— We then evaluate the proposed DLIF against ConvNeXt [liu2022convnet], a recent ConvNet architecture that has been recognized as a strong counterpart to Transformer models. We employ the default configuration of ConvNeXt but convert its 2D modules into 3D modules with the input remaining consistent with the aforementioned specifications.
  • EPICA—The final baseline method compared in this study is EPICA [naganawa2005extraction, naganawa2005omission], a statistical decomposition-based method discussed in the Introduction section. EPICA requires a scaling factor, typically derived from a single blood sample, to calibrate the estimated IDIF to align with the amplitude of the actual AIF. In our evaluation, we used the peak value of the ground-truth AIF as the scaling factor for the estimated IDIF.
@@ 4.4 Evaluation metrics
The DLIFs generated by the proposed framework have been applied in Logan graphical analysis. However, it is important to note that the potential applications of the proposed DLIF are not confined to this specific analysis technique. Logan graphical analysis employs linear regression to analyze the data after a specified time, estimating the slope of the resulting line to determine the total volume of distribution, as discussed in [carson2005tracer]. In this context, an exact match between the DLIF and the actual AIF before this chosen time is not critical. The crucial aspect is that the integral of the DLIF (i.e., the area under its curve) matches that of the AIF prior to this time. After this time point, the curves should align as closely as possible. This precise alignment is vital for accurately estimating the slope, which is crucial for the correct calculation of the total volume of distribution. Given these considerations, using metrics such as the mean squared error or the mean absolute error to assess the accuracy of the DLIFs may not provide the most meaningful insights. This is because such metrics could be disproportionately influenced by the errors in matching the peak values, which are significantly higher in amplitude compared to the input function values at later time points.
4.4.1 Pearson’s correlation coefficient ($\boldsymbol{r}$)
To evaluate the accuracy of AIFs estimated by various methods, we adopt Pearson’s correlation coefficient or $\boldsymbol{r}$, a metric also employed in related research on AIF estimation in perfusion CT [de2021aifnet]. This metric quantifies the correlation between the estimated AIF and the ground truth, and is mathematically defined as:
@@ 4.4.4 Logan graphical analysis
The Logan graphical analysis is widely used for quantifying dynamic PET images of a reversible tracer, where the tissue time-activity curve (TAC) is mathematically transformed and plotted against “normalized time” to estimate the total distribution volume ($V_{T}$). It is mathematically defined as:
$\frac{\int_{0}^{t}C_{T}(\tau)\,d\tau}{C_{T}(t)}=V_{T}\frac{\int_{0}^{t}C_{P}(\tau)\,d\tau}{C_{T}(t)}+\text{intercept}$ (12)
, where $C_{T}(t)$ is the measured tissue activity and $C_{p}(t)$ is the input curve. For the reversible compartment, this expression results in a straight line with a slope of $V_{T}$ after an equilibration time (t*). In previous studies, we have demonstrated that Logan graphical analysis with a t* = 30 minutes could provide robust $V_{T}$ values that agree with those produced from compartmental models [endres2009initial].
@@ 4.5 Ablation study and component analysis
  • Direct Est.”—This model directly predicts AIF values at specific discrete intervals, in accordance with the AIF values acquired using arterial sampling.
  • Exp.”—This model generates parameters for the exponential-sigmoid basis functions, which are then superimposed to create a continuous estimate of the AIF.
  • Exp. + Peak”—In addition to generating parameters for the exponential-sigmoid basis functions, this model adds a scaling factor to adjust the amplitude of the estimated AIF.
  • Exp. + Sparse”—In addition to generating parameters for the exponential-sigmoid basis functions, this model applies a sparsity constraint to encourage employing a smaller set of basis functions.
  • Exp. + Peak + Sparse”—This model not only generates parameters for the exponential-sigmoid basis functions and a scaling factor but also includes a sparsity constraint.
  • Gaussian”—This model outputs parameters for Gaussian basis functions, which are used to model the AIF as a sum of Gaussian curves.
  • Gaussian + Peak”—In addition to the Gaussian basis function parameters, this model includes a scaling factor to scale the amplitude of the estimated AIF.
  • Gaussian + Sparse”—In addition to the Gaussian basis function parameters, this model adds a sparsity constraint on the Gaussian basis functions, aiming to reduce the number of basis functions required for representing the AIF.
  • Gaussian + Peak + Sparse”—This model generates parameters for the Gaussian basis functions, includes a peak scaling factor and applies a sparsity constraint during training.
Fig. 6: Qualitative comparison of baseline versus proposed techniques using five-fold cross-validation on both genotypes, with each column representing results from a single patient. The first two columns present results from two patients with HAB, contrasted with the last two columns showcasing results from MAB patients. The first row depicts overall estimated input functions, and the second and the third rows offer zoom-in views within the 0 to 7 minute and $2$ to 90 minute timeframes, respectively.
@@ 5.1 Composition of DLIFs using basis functions
First, configurations employing Gaussian basis functions typically activate a larger number of functions compared to those using exponential-sigmoid basis functions. Even when applying a sparsity constraint, multiple Gaussian functions remain active (e.g., “Gaussian + Peak + Sparse”), enabling DLIFs to accurately capture subtle details of the ground-truth AIF. Conversely, exponential-sigmoid configurations often utilize fewer basis functions. Notably, under sparsity constraints, the DNN can converge to solutions involving a minimal number of exponential-sigmoid functions—such as the “Exp. + Peak + Sparse” configuration—where only one function remains active, effectively suppressing the others. This outcome highlights the intrinsic compatibility between the exponential-sigmoid shape and true AIF curves, facilitating concise yet accurate representations.
@@ 5.2 Ablation and component analysis
Next, we analyze the key elements that contribute to the improved quantitative performance of the proposed DLIF framework. To evaluate this, we used Pearson’s correlation coefficient ($\boldsymbol{r}$) to quantify the alignment of the DLIF shapes with the ground-truth AIFs. Additionally, we introduce a novel metric, Intersection-over-Union (IoU), which measures the overlap between the areas under the estimated and true input function curves. Unlike metrics focused on matching peak values, IoU emphasizes the integral areas, which are critical for downstream analyses such as Logan graphical analysis [logan1990graphical]. Details of these metrics are provided in the Methods section.
To refine the evaluation, we partition the input functions into two temporal segments: before and after 30 minutes, corresponding to the peak and stable phases of the input functions. Figure 5 presents the results of the component analysis using five-fold cross-validation on the validation datasets. The first plot in Fig. 5 depicts $\boldsymbol{r}$ values, representing the shape correspondence between the DLIFs and the actual AIFs. Although direct estimation of input functions (labeled “Direct Est.”) achieves an average $\boldsymbol{r}$ of 0.958, DLIF configurations using basis functions, either exponential or Gaussian, consistently outperform this approach. Specifically, the “Exp.”, “Gaussian”, “Gaussian+Peak”, and “Gaussian+Peak+Sparse” configurations attain average $\boldsymbol{r}$ values of 0.965, 0.964, 0.961, and 0.964, respectively. The second plot evaluates IoU, capturing the agreement in amplitudes and the total areas under the curves. Here, the “Exp.”, “Gaussian”, “Gaussian+Peak”, and “Gaussian+Peak+Sparse” configurations achieve mean IoU scores of 0.787, 0.777, 0.781, and 0.780, respectively, slightly outperforming “Direct Est.” (0.776).
Performance in the different temporal phases reveals further insight. During the peak phase (<30 minutes), the top-performing configurations, “Exp.”, “Gaussian+Peak”, and “Gaussian+Peak+Sparse”, record average IoU scores of 0.786, 0.781, and 0.779, respectively, compared to 0.776 for “Direct Est.”. During the tail phase (>30 minutes), these configurations achieve average IoU scores of 0.808, 0.803, and 0.799, respectively, compared to 0.780 for “Direct Est.”. Notably, the “Gaussian” configuration slightly underperforms in the peak phase, with an average IoU of 0.776.
Furthermore, the scatter plots demonstrate tighter distributions with shorter tails for “Exp.”, “Gaussian+Peak”, and “Gaussian+Peak+Sparse” compared to “Direct Est.”, indicating more consistent performance. Based on these findings, we selected the top three configurations—“Exp.”, “Gaussian+Peak”, and “Gaussian+Peak+Sparse”—for further quantitative and qualitative comparisons against baseline methods, as detailed in the subsequent section.
Table 1: Quantitative analysis of different methods on test sets not exposed during training. The results are based on five-fold cross-validation. Note that the time intervals “30-” and “30+” refer to the periods of 0-30 minutes and 60-90 minutes, respectively. The top-performing scores are highlighted in bold. The symbol “*” denotes a statistically significant improvement, with a $p$-value < 0.05, as determined by a Wilcoxon signed-rank test with Bonferroni adjustment for multiple comparisons.
@@
[…]